# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 6 $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
We define the invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n^{2}$ (which we think of as an adjacency matrix for a graph with $n$ vertices) and randomness of exactly $\log\left(n\right)$ bits (formally, we use $\lceil\log\left(n\right)\rceil$ bits). We will later define the circuits $C^{0}$, $C^{1}$, as a function of the circuit sequence $C$.
@@
      • In case $k_{i,j}\leq i$ the circuit $C_{\left(i,j\right),\rightarrow}$ computes the graph $\sigma_{\left(i,k_{i,j}\right)}\left(G\right)$ as its output, where $\sigma_{\left(i,k_{i,j}\right)}\in S_{i}$ is the permutation on the elements $[i]$, that acts trivially on all elements that are in $[i]\setminus\{i,k_{i,j}\}$, and swaps between $i$ and $k_{i,j}$.
We defined what are the forward computations, and the definition of the inverse circuits follows: For every $i\in\{n,n-1,\cdots,3,2\}$, $j\in[n]$, the circuit $C_{\left(i,j\right),\leftarrow}$ is exactly identical to the circuit $C_{\left(i,j\right),\rightarrow}$. It can be verified by the reader that indeed for every $i\in\{n,n-1,\cdots,3,2\}$, $j\in[n]$, the circuits $C_{\left(i,j\right),\rightarrow}$, $C_{\left(i,j\right),\leftarrow}$ are inverses of each other, because making the same swap twice leads to the identity permutation.
@@
The next step of our proof is to show that the reduction is correct, that is, that when $\left(G_{0},G_{1}\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits, denoted $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ respectively, are statistically close, and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far.
Both soundness and completeness follow rather easily at this point. For soundness, in case $\left(G_{0},G_{1}\right)\in\prod_{\text{NO}}$, it means the graphs are non-isomorphic, which, by Fact 3.1, means that the supports of the distributions $\sigma\left(G_{0}\right)$, $\sigma\left(G_{1}\right)$ are disjoint, and thus $\left\lVert C^{0}-C^{1}\right\rVert_{TV}\geq 1-\frac{n}{2^{n}}$. For completeness, in case $\left(G_{0},G_{1}\right)\in\prod_{\text{YES}}$, it means the graphs are isomorphic, which, by Fact 3.1, means that the distributions $\sigma\left(G_{0}\right)$, $\sigma\left(G_{1}\right)$ are identical, and thus $\left\lVert C^{0}-C^{1}\right\rVert_{TV}\leq\frac{n}{2^{n}}$. ∎
@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$\mathsf{GapCVP}_{c\cdot n^{2}}\leq_{p}\mathsf{SISD}_{a(n),b(n),r(n)}\enspace,$
@@
Let $\left(\mathbf{B},\mathbf{t},d\right)$ an input for the $\mathsf{GapCVP}_{g(n)}$ problem, for $g(n)=c_{g}\cdot n^{2}$ for $c_{g}:=256\cdot\frac{c_{0}}{c_{1}}\cdot\pi$, where $c_{0},c_{1}\in\mathbb{R}_{>0}$ are two positive absolute constants, specified later, and $n\in\mathbb{N}$ is the dimension of the lattice basis $\mathbf{B}$ and the vector $\mathbf{t}$, and $b\in\mathbb{N}$ is the size of the binary representation of the largest (in absolute value) number in $\mathbf{B}$ and the vector $\mathbf{t}$. So, the description size of the input is
@@
The reduction executes in classical deterministic polynomial time and will output a pair $\left(C^{0},C^{1}\right)$ of invertible circuit sequences (as in Definition 4.6). Let $\ell:=\ell_{\mathcal{L}}+1+\ell_{G}+1$, where $\ell_{\mathcal{L}}:=\left(m+1\right)\cdot n$ and $\ell_{G}:=\kappa\cdot\left(\beta+1\right)\cdot n$, such that,
@@
Definition of the circuit sequence $C$ and the pair $\left(C^{0},C^{1}\right)$.
We define the invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n\cdot\left(s+1\right)$ (which we think of as an integer vector $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$) and randomness bounded by $1$ (that is, each circuit either uses a single bit of randomness or it is deterministic, and does not use randomness at all). In the below circuits, all arithmetic (which comes down to additions and multiplications) is modulo in absolute value of $12^{s}-$ 1.
@@
    Constructing a random lattice vector with uniform coordinates vector in ($\mathbb{Z}^{n}_{\left(0,2^{m+1}-1\right)}$. For $i\in[m+1]$, $j\in[n]$, the circuit $C_{\left(1,i,j\right),\rightarrow}$, given main input $\mathbf{v}$ and auxiliary input randomness $z_{1,i,j}\in\{0,1\}$, adds the vector $z_{1,i,j}\cdot 2^{i-1}\cdot\mathbf{B}_{j}$, where $\mathbf{B}_{j}$ is the $j$-th column of the basis $\mathbf{B}$. Note that this step uses $\ell_{\mathcal{L}}:=\left(m+1\right)\cdot n$ circuits, where each uses a single bit of randomness.
@@
    Stabilizing the coordinate vectors to ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, by subtraction. The circuit $C_{\left(2,1\right),\rightarrow}$, which is deterministic and does not use randomness, given input $\mathbf{v}$, outputs $\mathbf{v}-2^{m}\cdot\sum_{j\in[n]}\mathbf{B}_{j}$. This step uses a single deterministic circuit.
@@
    Adding approximated Gaussian noise. For $i_{1}\in[\kappa],i_{2}\in[\beta+1],i_{3}\in[n]$, the circuit $C_{\left(3,i_{1},i_{2},i_{3}\right),\rightarrow}$, given input $\mathbf{v}$ and randomness $z_{\left(3,i_{1},i_{2},i_{3}\right)}\in\{0,1\}$, adds the vector
@@
    where for $j\in[n]$, the vector $e_{j}$ is the $j$-th standard basis vector, i.e., $e_{j}:=\left(0_{1},\cdots,0_{j-1},1_{j},0_{j+1},\cdots,0_{n}\right)^{T}$. Note that this step uses $\ell_{G}:=\kappa\cdot\left(\beta+1\right)\cdot n$ circuits, where each uses a single bit of randomness.
@@
    Stabilizing the Gaussian noise by subtraction. The circuit $C_{\left(4,1\right),\rightarrow}$, which is deterministic and does not use randomness, given input $\mathbf{v}$, outputs
    $\mathbf{v}-\left(\frac{\kappa}{2}\cdot\left(2\cdot 2^{\beta}-1\right)\left(1_{1},1_{2},\cdots,1_{n}\right)^{T}\right)\enspace.$
@@
    $\left(\frac{\kappa}{2}\cdot\left(2\cdot 2^{\beta}-1\right)\left(1_{1},1_{2},\cdots,1_{n}\right)^{T}\right)\enspace,$
@@
  • The next step 2 of the algorithm is intended to move the distribution to be uniform with expectation $\approx 0$, and to this end the circuit $C_{\ell_{\mathcal{L}}+1,\rightarrow}$ subtracts $2^{m}\cdot\sum_{j\in[n]}\mathbf{B}_{j}$. At the end of this step the output distribution is accordingly $\sum_{j\in[n]}\left(M_{j}-2^{m}\right)\cdot\mathbf{B}_{j}$, where for every $j\in[n]$, the variables $M_{j}-2^{m}$ are i.i.d. samples from the uniform distribution over $\{-2^{m},\cdots,-1,0,1,\cdots,2^{m}-1\}$. In other words, at the end of step 2 we have a random lattice vector, over a uniformly random coordinates vector in the restricted set ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$.
@@
  • The next step is intended to stabilize as much as possible the Gaussian noise to center around 00. To this end, step 4 subtracts $\frac{\kappa}{2}\left(2^{\beta}+2^{\beta}-1\right)$ ($\kappa/2$ is an integer) from each of the $n$ coordinates of $\mathbf{v}$, this can also be thought of as subtracting $\frac{\kappa}{2}\cdot 2^{\beta}$ followed by subtracting $\frac{\kappa}{2}\cdot\left(2^{\beta}-1\right)$.
@@
  • The output distribution of $C^{0}$ is the variable $\left(\mathbf{v}+\mathbf{e}\right)\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right),2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right)}$, where $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n\right),2^{m+b}\cdot n\right)}$ is a random lattice vector with a uniformly random coordinates vector $\mathbf{a}\in\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, and $\mathbf{e}\in\mathbb{Z}^{n}_{\left(-\kappa\cdot 2^{\beta},\kappa\cdot 2^{\beta}\right)}$ such that for each $j\in[n]$, the coordinate $\mathbf{e}_{j}$ is a random variable that’s the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$.
@@
The next step of our proof is to show that the reduction is correct, that is, that when $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits, denoted $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ respectively, are statistically close, and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far.
@@
Completeness of the reduction, first part.
@@
In our next step we would like to use the fact that adding a lattice vector to a uniform distribution over that lattice, does not change the distribution. However, the lattice is infinite and we cannot sample a uniformly random vector over the entire lattice. Formally, since $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, thus there exist $\mathbf{s}\in\mathcal{L}_{\mathbf{B}}$ and $\mathbf{e}_{0}\in\mathbb{Z}^{n}$ such that $\mathbf{t}=\mathbf{s}+\mathbf{e}_{0}$ and $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$. It follows that,
@@
We use the fact that we took the bound $M$ (which is effectively the sample space of our coordinates for the vector $\mathbf{v}$) to be exponentially (in the input dimension $n$) larger than the vectors in $\mathbf{B}$ and $\mathbf{t}$. Formally, note that since the norm of $\mathbf{s}$ is $\leq\left\lVert\mathbf{t}\right\rVert+\left\lVert\mathbf{e}_{0}\right\rVert\leq\left\lVert\mathbf{t}\right\rVert+d$, then $\left\lVert\mathbf{s}\right\rVert\cdot 2^{n}\leq M$. It follows that the distribution $\mathbf{v}-\mathbf{s}$ has total variation distance bounded by $2^{-n}$ to $\mathbf{v}$, and the above sum is bounded by
@@
where $\tilde{B}:=\lceil\kappa\cdot\tilde{\sigma}\rceil$, $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$, and the distribution $D^{n}_{\tilde{B}}$ is the $n$-dimensional discrete truncated Gaussian distribution, as defined in Definition 3.12. We use once again the fact that adding a numerical random variable to a pair of sums does not make it easier to distinguish the sums, and thus $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}\leq\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$, and overall we have,

### v2

@@ 6 $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
We define an invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n^{2}$ (which we think of as an adjacency matrix for a graph with $n$ vertices) and randomness of exactly $\log\left(n\right)$ bits (formally, we use $\lceil\log\left(n\right)\rceil$ bits). We will later define the circuits $C^{0}$, $C^{1}$, as a function of the circuit sequence $C$.
@@
      • In case $k_{i,j}\leq i$ the fixed-randomness circuit $C_{\left(i,j\right),\rightarrow}\left(\cdot;z_{i,j}\right)$ computes the graph $\sigma_{\left(i,k_{i,j}\right)}\left(G\right)$ as its output, where $\sigma_{\left(i,k_{i,j}\right)}\in S_{i}$ is a swap between elements $i$ and $k_{i,j}$, that is, it is a permutation on the elements $[i]$, that acts trivially on all elements that are in $[i]\setminus\{i,k_{i,j}\}$, and swaps between $i$ and $k_{i,j}$.
We defined what are the forward computations, and the definition of the inverse circuits follows: For every $i\in\{n,n-1,\cdots,3,2\}$, $j\in[n]$, the circuit $C_{\left(i,j\right),\leftarrow}$ is exactly identical to the circuit $C_{\left(i,j\right),\rightarrow}$. It can be verified by the reader that indeed for every $i\in\{n,n-1,\cdots,3,2\}$, $j\in[n]$, the circuits $C_{\left(i,j\right),\rightarrow}$, $C_{\left(i,j\right),\leftarrow}$ are inverses of each other for every fixed randomness $z_{i,j}\in\{0,1\}^{\lceil\log(n)\rceil}$, because making the same swap twice leads to the identity permutation.
@@
The last step of our proof is to show that the reduction is correct, that is, that when $\left(G_{0},G_{1}\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits, denoted $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ respectively, are statistically close, and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far.
Both soundness and completeness follow rather easily at this point. For soundness, in case $\left(G_{0},G_{1}\right)\in\prod_{\text{NO}}$, it means the graphs are non-isomorphic, which, by Fact 3.1, means that the supports of the distributions $\sigma\left(G_{0}\right)$, $\sigma\left(G_{1}\right)$ are disjoint, and thus $\left\lVert C^{0}-C^{1}\right\rVert_{TV}\geq 1-\frac{n}{2^{n}}$. For completeness, in case $\left(G_{0},G_{1}\right)\in\prod_{\text{YES}}$, it means the graphs are isomorphic, which again by Fact 3.1, means that the distributions $\sigma\left(G_{0}\right)$, $\sigma\left(G_{1}\right)$ are identical, and thus $\left\lVert C^{0}-C^{1}\right\rVert_{TV}\leq\frac{n}{2^{n}}$. ∎
@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$\mathsf{GapCVP}_{c\cdot n^{2}}\leq_{p}\mathsf{SISD}_{\left(a(n),b(n),r(n)\right)}\enspace,$
@@
Let $\left(\mathbf{B},\mathbf{t},d\right)$ an input for the $\mathsf{GapCVP}_{g(n)}$ problem, for $g(n)=c_{g}\cdot n^{2}$ where $c_{g}:=256\cdot\frac{c_{0}}{c_{1}}\cdot\pi$, where $c_{0},c_{1}\in\mathbb{R}_{>0}$ are two positive absolute constants we will specify later, $n\in\mathbb{N}$ is the dimension of the lattice basis $\mathbf{B}$ and the vector $\mathbf{t}$, and $b\in\mathbb{N}$ is the size of the binary representation of the largest (in absolute value) number in $\mathbf{B}$ and the vector $\mathbf{t}$. So, the description size of the input is
@@
The reduction executes in classical deterministic polynomial time and will output a pair $\left(C^{0},C^{1}\right)$ of $1$-invertible circuit sequences (as in Definition 4.8). Let $\ell:=\ell_{\mathcal{L}}+1+\ell_{G}+1$, where $\ell_{\mathcal{L}}:=\left(m+1\right)\cdot n$ and $\ell_{G}:=\kappa\cdot\left(\beta+1\right)\cdot n$, such that,
@@
Definition of the circuit sequence $C$ and an pair $\left(C^{0},C^{1}\right)$.
We define an invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n\cdot\left(s+1\right)$ (which we think of as an integer vector $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$) and randomness bounded by $1$ (that is, each circuit either uses a single bit of randomness or it is deterministic, and does not use randomness at all). In the below circuits, all arithmetic (which comes down to additions and multiplications) is modulo in absolute value of $12^{s}-$ 1.
@@
    First $\ell_{\mathcal{L}}$ circuits: Constructing a random lattice vector with a uniform coordinates vector in ($\mathbb{Z}^{n}_{\left(0,2^{m+1}-1\right)}$. For $i\in[m+1]$, $j\in[n]$, the circuit $C_{\left(1,i,j\right),\rightarrow}$, given main input $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$ and auxiliary input randomness $z_{1,i,j}\in\{0,1\}$, adds the vector $z_{1,i,j}\cdot 2^{i-1}\cdot\mathbf{B}_{j}$, where $\mathbf{B}_{j}$ is the $j$-th column of the basis $\mathbf{B}$.
@@
    Circuit $\ell_{\mathcal{L}}+1$: Stabilizing the coordinate vectors to ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, by subtraction. The circuit $C_{\left(2,1\right),\rightarrow}$, which is deterministic and does not use randomness, given input $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$, outputs $\mathbf{v}-2^{m}\cdot\sum_{j\in[n]}\mathbf{B}_{j}$. This step uses a single deterministic circuit.
@@
    Circuits $\ell_{\mathcal{L}}+1+i$ for $i\in[\ell_{G}]$: Adding approximated Gaussian noise. For $i_{1}\in[\kappa],i_{2}\in[\beta+1],i_{3}\in[n]$, the circuit $C_{\left(3,i_{1},i_{2},i_{3}\right),\rightarrow}$, given input $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$ and randomness $z_{\left(3,i_{1},i_{2},i_{3}\right)}\in\{0,1\}$, adds the vector
@@
    where for $j\in[n]$, the vector $e_{j}$ is the $j$-th standard basis vector, i.e., $e_{j}:=\left(0_{1},\cdots,0_{j-1},1_{j},0_{j+1},\cdots,0_{n}\right)^{T}$.
@@
    Circuit $\ell_{\mathcal{L}}+1+\ell_{G}+1$: Stabilizing the Gaussian noise by subtraction. The circuit $C_{\left(4,1\right),\rightarrow}$, which is deterministic and does not use randomness, given input $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$, outputs
    $\mathbf{v}-\left(\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)\left(1_{1},1_{2},\cdots,1_{n}\right)^{T}\right)\enspace.$
@@
    $\left(\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)\left(1_{1},1_{2},\cdots,1_{n}\right)^{T}\right)\enspace,$
@@
  • The next step 2 of the algorithm is intended to move the distribution of the coordinates vector $\left(M_{1},\cdots,M_{n}\right)$ be uniform with expectation $\approx 0$, and to this end the circuit $C_{\ell_{\mathcal{L}}+1,\rightarrow}$ subtracts $2^{m}\cdot\sum_{j\in[n]}\mathbf{B}_{j}$. At the end of this step the output distribution is accordingly $\sum_{j\in[n]}\left(M_{j}-2^{m}\right)\cdot\mathbf{B}_{j}$, where for every $j\in[n]$, the variables $M_{j}-2^{m}$ are i.i.d. samples from the uniform distribution over $\{-2^{m},\cdots,-1,0,1,\cdots,2^{m}-1\}$. In other words, at the end of step 2 we have a random lattice vector, over a uniformly random coordinates vector in the restricted set ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$.
@@
  • The next step is intended to stabilize the Gaussian noise to center around 00 (in expectation). To this end, step 4 subtracts $\frac{\kappa}{2}\left(2^{\beta+1}-1\right)$ ($\kappa/2$ is an integer) from each of the $n$ coordinates of $\mathbf{v}$. Since this is exactly the expectation of the random variable we add to each coordinate, the average of Gaussian noise becomes zero.
@@
  • The output distribution of $C^{0}$ is the variable $\left(\mathbf{v}+\mathbf{e}\right)\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right),2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right)}$, where $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n\right),2^{m+b}\cdot n\right)}$ is a random lattice vector with a uniformly random coordinates vector $\mathbf{a}\in\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, and $\mathbf{e}\in\mathbb{Z}^{n}_{\left(-\kappa\cdot 2^{\beta},\kappa\cdot 2^{\beta}\right)}$ such that for each $j\in[n]$, the coordinate $\mathbf{e}_{j}$ is a random variable defined as follows. Since we subtracted at the end $\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)$, it is equivalent to subtracting $2^{\beta}$, $\frac{\kappa}{2}$ times followed by subtracting $\left(2^{\beta}-1\right)$, $\frac{\kappa}{2}$ times. We can thus consider the variable $\mathbf{e}_{j}$ as the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$.
@@
We show that the reduction is correct, that is, that when $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits, denoted $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ respectively, are statistically close, and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far.
@@
Completeness of the reduction, first part (reducing the analysis to analyzing total variation distance of the noise vector $\mathbf{e}$).
@@
We would like to use the fact that adding a lattice vector to a uniform distribution over that lattice, does not change the distribution. However, the lattice is infinite and we cannot sample a uniformly random vector over the entire lattice. Formally, since $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, thus there exist $\mathbf{s}\in\mathcal{L}_{\mathbf{B}}$ and $\mathbf{e}_{0}\in\mathbb{Z}^{n}$ such that $\mathbf{t}=\mathbf{s}+\mathbf{e}_{0}$ and $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$. It follows that,
@@
We use the fact that we took the bound $M$ (which is effectively the sample space of our coordinates for the vector $\mathbf{v}$) to be exponentially (in the input dimension $n$) larger than the vectors in $\mathbf{B}$ and $\mathbf{t}$. Formally, note that since the norm of $\mathbf{s}$ is $\leq\left\lVert\mathbf{t}\right\rVert+\left\lVert\mathbf{e}_{0}\right\rVert\leq\left\lVert\mathbf{t}\right\rVert+d$, then $\left\lVert\mathbf{s}\right\rVert\cdot 2^{n}\leq M$. Since the variable $\mathbf{v}$ has a uniformly random coordinates vector in ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, the length $\mathbf{s}$ is tiny compared to $\mathbf{v}$ with overwhelming probability. It follows that the distribution $\mathbf{v}-\mathbf{s}$ has total variation distance bounded by $2^{-n}$ to $\mathbf{v}$, and the above sum of total variation distances is bounded by
@@
where $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$, $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$, and the distribution $D^{n}_{\tilde{B}}$ is the $n$-dimensional discrete truncated Gaussian distribution, as defined in Definition 3.12. We use once again the fact that adding a numerical random variable to a pair of sums does not make it easier to distinguish the sums, and thus $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}\leq\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$, and overall we have,